# On finite group scheme-theoretical categories, II

### v1

@@ 6.1. The étale case
Assume $G$ is étale, i.e, $G=G(k)$. The following result is known in characteristic 00 [GN, O].
@@
    For any $(H,H)$-double coset $Z$ with representative $g\in G$, and simple $V\in{\rm Corep}_{k}(H^{g},\xi_{g}^{-1})$, we have
@@
    ${\mathscr{C}}\cong\bigoplus_{Z\in Y}{\rm Corep}_{k}(H^{g},\xi_{g}^{-1}),$
@@ 6.2. The connected case
For any $V\in{\rm Corep}_{k}(H)$, set ${\rm F}_{V}=\left({\rm id}\otimes\alpha\right)_{21}$ (see §3.4). We have
@@
For any $V\in{\rm Corep}_{k}(H)$, the following hold:
@@
    $\mathbf{F}:=\mathbf{F}_{H}:{\rm Corep}_{k}(H)\xrightarrow{\cong}{\mathscr{C}}_{H},\,\,\,V\mapsto\iota_{*}\left(\mathscr{O}(H)\otimes_{k}V,\lambda^{1}_{V},\rho^{1}_{V}\right),$
@@
    For any simple $V\in{\rm Corep}_{k}(H)$, we have
@@
    For any $V\in{\rm Corep}_{k}(H)$, $\text{\rm FPdim}\left(P_{{\mathscr{C}}}\left(\mathbf{F}\left(V\right)\right)\right)=\frac{|G|}{|H|}\dim\left(P_{H}\left(V\right)\right)$.
@@
Let $\mathfrak{g}$ be a finite dimensional restricted $p$-Lie algebra, $\mathfrak{h}\subset\mathfrak{g}$ a restricted $p$-Lie subalgebra, and $G,H$ the associated finite group schemes. Consider the finite tensor category ${\mathscr{C}}:={\mathscr{C}}\left(G,\omega,H,\psi\right)$. (See [G3] for examples of nontrivial cocycles.) By Theorem 6.6, we have ${\mathscr{C}}=\overline{{\rm Corep}_{k}(\mathfrak{h})}$.
@@
is the free $\mathscr{O}(G)$-module of rank $|H|$ with trivial left $\mathscr{O}(H)$-cocation and right $\mathscr{O}(H)$-cocation ${\rm id}\otimes\Delta_{\psi}$. So, ${\mathscr{C}}\cong{\rm Coh}(G,\omega)$ as abelian categories, but not necessarily as tensor categories (see [GS, Example 6.8]). ∎
@@ 6.3. The normal case
Assume $H$ is normal in $G$. Then $G/H=H\backslash G$ is a finite group scheme, and the quotient morphism $\pi:G\to G/H$ (2.15) is a group scheme morphism.
@@
    ${\rm Coh}(G/H)\boxtimes{\rm Corep}_{k}(H)\xrightarrow{\cong}{\mathscr{C}},\,\,\,U\boxtimes V\mapsto\pi^{*}U\otimes\mathbf{F}(V).$
@@
$\pi^{*}\delta_{\bar{g}}\otimes-:\overline{{\mathscr{C}}_{H}}\xrightarrow{\cong}\overline{{\mathscr{C}}_{gH}},\,\,\,S\mapsto\pi^{*}\delta_{\bar{g}}\otimes S.$
@@
${\rm Coh}(G/H)_{\bar{1}}\boxtimes{\rm Corep}_{k}(H)\xrightarrow{\cong}\overline{{\mathscr{C}}_{H}},\,\,\,U\boxtimes V\mapsto\pi^{*}U\otimes\mathbf{F}(V),$
@@
${\rm Coh}(G/H)_{\bar{g}}\boxtimes{\rm Corep}_{k}(H)\xrightarrow{\cong}\overline{{\mathscr{C}}_{gH}},\,\,\,U\boxtimes V\mapsto\pi^{*}U\otimes\mathbf{F}(V),$
@@ 7. The center of ${\rm Coh}(G,\omega)$
Throughout this section, we fix $(G,\omega)$ 2.7. Let $\Delta\colon G\to G\times G$ be the diagonal map, $\mathbb{G}:=G\times G$, and $\mathbb{H}:=\Delta(G)$. Let $\mathscr{Z}(G,\omega):=\mathscr{Z}({\rm Coh}(G,\omega))$ be the center of ${\rm Coh}(G,\omega)$. Recall that objects of $\mathscr{Z}(G,\omega)$ are pairs $(X,{\rm c})$, where $X\in{\rm Coh}(G,\omega)$ and
@@
(7.1) $\mathscr{Z}(G,\omega)\xrightarrow{\cong}\left({\rm Coh}(G,\omega)\boxtimes{\rm Coh}\left(G,\omega\right)^{{}^{\rm rev}}\right)^{*}_{{\rm Coh}(G,\omega)}$
@@
Let $p_{1},p_{2}:\mathbb{G}\twoheadrightarrow G$ be the projection morphisms, and let
@@
(7.2) ${\rm Coh}(\mathbb{G},\widetilde{\omega})^{*}_{{\mathscr{M}}(\mathbb{H},1)}\xrightarrow{\cong}\left({\rm Coh}(G,\omega)\boxtimes{\rm Coh}\left(G,\omega\right)^{{}^{\rm rev}}\right)^{*}_{{\rm Coh}(G,\omega)}.$
@@
(7.3) ${\mathscr{C}}\left(\mathbb{G},\widetilde{\omega},\mathbb{H},1\right)\cong\mathscr{Z}(G,\omega).$
@@
(7.4) ${\rm Coh}^{(G)}(G,\omega)\cong\mathscr{Z}(G,\omega),$
@@
(7.5) ${\mathscr{C}}\left(\mathbb{G},\widetilde{\omega},\mathbb{H},1\right)\cong{\rm Coh}^{(G)}(G,\omega).$
@@ 7.1. The structure of ${\mathscr{C}}\left(\mathbb{G},\tilde{\omega},\mathbb{H},1\right)$
    $\mathbf{F}_{{\rm Z}}:{\rm Corep}_{k}(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})\xrightarrow{\cong}\mathscr{C}_{{\rm Z}},\,\,\,V\mapsto\iota_{{\rm Z}*}\left(\mathscr{O}\left({\rm Z}\right)\otimes_{k}V,\lambda_{V}^{{\rm g}},\rho^{{\rm g}}_{V}\right).$
@@
    $\mathbf{F}_{\mathbb{H}}:{\rm Corep}_{k}(\mathbb{H})\xrightarrow{\cong}\mathscr{C}_{\mathbb{H}},\,\,\,V\mapsto\iota_{\mathbb{H}*}\left(\mathscr{O}\left(\mathbb{H}\right)\otimes_{k}V,\lambda_{V}^{1},\rho^{1}_{V}\right).$
@@
    There is a bijection between equivalence classes of pairs $({\rm Z},V)$, where ${\rm Z}\in{\rm Y}(k)$ is a closed point with representative ${\rm g}\in{\rm Z}(k)$, and $V\in{\rm Corep}_{k}(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})$ is simple, and simple objects of ${\mathscr{C}}$, assigning $({\rm Z},V)$ to $\mathbf{F}_{{\rm Z}}(V)$. Moreover, we have a direct sum decomposition of abelian categories
@@
    For any $V\in{\rm Corep}_{k}(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})$, we have $\mathbf{F}_{{\rm Z}}(V)^{*}\cong\mathbf{F}_{{\rm Z}^{-1}}(V^{*})$.
@@
    For any ${\rm Z}\in{\rm Y}(k)$ with representative ${\rm g}\in{\rm Z}(k)$, and $V$ in ${\rm Corep}_{k}(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})$, we have
@@
    For any ${\rm Z}\in{\rm Y}(k)$ with representative ${\rm g}\in{\rm Z}(k)$, and simple $V\in{\rm Corep}_{k}(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})$, we have
@@
Equivalence classes of fiber functors on ${\mathscr{C}}\left(\mathbb{G},\widetilde{\omega},\mathbb{H},1\right)$ are classified by equivalence classes of pairs $(\mathbb{K},\eta)$, where $\mathbb{K}\subset\mathbb{G}$ is a closed subgroup scheme and $\eta\in C^{2}(\mathbb{K},\mathbb{G}_{m})$, such that $d\eta=\iota_{\mathbb{K}}^{\sharp\otimes 3}(\widetilde{\omega})$, $\mathbb{K}\mathbb{H}=\mathbb{G}$, and $\xi_{1}^{-1}\in Z^{2}(\mathbb{K}\cap\mathbb{H},\mathbb{G}_{m})$ is nondegenerate.
@@ 7.2. The structure of ${\rm Coh}^{(G)}(G,\omega)$
(7.6) $\alpha_{g}:\mathscr{O}(G)\twoheadrightarrow\mathscr{O}(G_{C}\backslash G),\,\,\,f\mapsto f_{1}\mathfrak{c}_{g}^{-1}\left(\iota_{g}^{\sharp}\left(f_{2}\right)\right)$
@@
    $\mathbf{F}_{C}:{\rm Rep}_{k}(G_{C},\omega_{g})\xrightarrow{\cong}\mathscr{Z}_{C},\,\,\,V\mapsto\iota_{C*}\left(\mathscr{O}(C)\otimes_{k}V,\rho_{V}^{g}\right),$
@@
    (Here, $\mathfrak{j}_{g}:G_{C}\backslash G\xrightarrow{\cong}C$ is the canonical scheme isomorphism.)
    In particular, $\mathbf{F}_{1}:{\rm Rep}_{k}(G)\xrightarrow{\cong}\mathscr{Z}_{1}\hookrightarrow\mathscr{Z}$ coincides with the canonical embedding of braided tensor categories.
@@ 7.4. $\mathscr{Z}(G,\omega)$ as $G(k)$-equivariantization
Set $\mathscr{D}:={\rm Coh}(G,\omega)$, and $\mathscr{D}^{\circ}:={\rm Coh}(G^{\circ},\omega^{\circ})$. By (2.4), we have
$\mathscr{D}\cong\mathscr{D}^{\circ}\rtimes G(k)\cong\bigoplus_{a\in G(k)}\mathscr{D}^{\circ}\boxtimes a$
as tensor categories, where the associativity constraint on the right hand side category is given by $\omega$ in the obvious way.
@@
Note that $\mathscr{Z}_{\mathscr{D}^{\circ}}(\mathscr{D})\cong{\rm Coh}^{(G^{\circ})}(G,\omega)$ as tensor categories, and the obvious forgetful tensor functor
@@
(7.11) $F:\mathscr{Z}(G,\omega)\xrightarrow{\cong}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right)^{G(k)}=\left(\bigoplus_{a\in G(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)\right)^{G(k)}.$
@@
(7.12) $F:\mathscr{Z}(G,\omega)\xrightarrow{\cong}\bigoplus_{C\in{\rm C}(k)}\mathscr{E}_{C}.$
@@
$F_{C}:\overline{\mathscr{Z}(G,\omega)_{C}}\xrightarrow{\cong}\mathscr{E}_{C}=\bigoplus_{a\in C(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)^{G(k)}.$
@@
$F_{1}:\overline{{\rm Rep}(G)}\xrightarrow{\cong}\mathscr{Z}\left(G^{\circ},\omega^{\circ}\right)^{G(k)}.$
@@
So, let ${\rm V}\in\mathscr{Z}(G,\omega)_{C}$ be simple. By Theorem 7.4, there exists a unique simple $V$ in ${\rm Rep}(G_{C},\omega_{g})$ such that ${\rm V}={\bf F}_{C}(V)=\mathscr{O}(C)\otimes_{k}V$. It follows that the forgetful image of $F({\rm V})$ in $\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right)$ (7.10) lies in $\bigoplus_{a\in C(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)$, so by the proof of [GNN, Theorem 3.5], $F({\rm V})$ is supported on $C$, as claimed. ∎

### v2

@@ 6.1. The étale case
Assume $G$ is étale, i.e, $G=G(k)$. The following result is known in characteristic 0 [GN, O].
@@
    For any $(H,H)$-double coset $Z$ with representative $g\in G$, and simple $V\in{\rm Rep}(H^{g},\xi_{g}^{-1})_{k}$, we have
@@
    ${\mathscr{C}}\simeq\bigoplus_{Z\in Y}{\rm Rep}(H^{g},\xi_{g}^{-1})_{k},$
@@ 6.2. The connected case
For any $V\in{\rm Corep}_{k}(\mathscr{O}(H))$, set ${\rm F}_{V}=\left({\rm id}\otimes\alpha\right)_{21}$ (see §3.4). We have
@@
For any $V\in{\rm Corep}_{k}(\mathscr{O}(H))$, the following hold:
@@
    $\mathbf{F}:=\mathbf{F}_{H}:{\rm Corep}_{k}(\mathscr{O}(H))\xrightarrow{\simeq}{\mathscr{C}}_{H},\,\,\,V\mapsto\iota_{*}\left(\mathscr{O}(H)\otimes_{k}V,\lambda^{1}_{V},\rho^{1}_{V}\right),$
@@
    For any simple $V\in{\rm Corep}_{k}(\mathscr{O}(H))$, we have
@@
    For any $V\in{\rm Corep}_{k}(\mathscr{O}(H))$, $\text{\rm FPdim}\left(P_{{\mathscr{C}}}\left(\mathbf{F}\left(V\right)\right)\right)=\frac{|G|}{|H|}\dim\left(P_{H}\left(V\right)\right)$.
@@
Let $\mathfrak{g}$ be a finite dimensional restricted $p$-Lie algebra, $\mathfrak{h}\subset\mathfrak{g}$ a restricted $p$-Lie subalgebra, and $G,H$ the associated finite group schemes. Consider the finite tensor category ${\mathscr{C}}:={\mathscr{C}}\left(G,\omega,H,\psi\right)$. (See [G3] for examples of nontrivial cocycles.) By Theorem 6.6, we have ${\mathscr{C}}=\overline{{\rm Rep}(\mathfrak{h})_{k}}$.
@@
is the free $\mathscr{O}(G)$-module of rank $|H|$ with trivial left $\mathscr{O}(H)$-cocation and right $\mathscr{O}(H)$-cocation ${\rm id}\otimes\Delta_{\psi}$. So, ${\mathscr{C}}\simeq{\rm Coh}(G,\omega)$ as abelian categories, but not necessarily as tensor categories (see [GS, Example 6.8]). ∎
@@ 6.3. The normal case
Assume $H$ is normal in $G$. Then $\GG/H=H\backslash G$ is a finite group scheme, and the quotient morphism $\pi:G\to G/H$ (2.17) is a group scheme morphism.
@@
    ${\rm Coh}(G/H)\boxtimes{\rm Rep}(H)_{k}\xrightarrow{\simeq}{\mathscr{C}},\,\,\,U\boxtimes V\mapsto\pi^{*}U\otimes\mathbf{F}(V).$
@@
$\pi^{*}\delta_{\bar{g}}\otimes-:\overline{{\mathscr{C}}_{H}}\xrightarrow{\simeq}\overline{{\mathscr{C}}_{gH}},\,\,\,S\mapsto\pi^{*}\delta_{\bar{g}}\otimes S.$
@@
${\rm Coh}(G/H)_{\bar{1}}\boxtimes{\rm Rep}(H)_{k}\xrightarrow{\simeq}\overline{{\mathscr{C}}_{H}},\,\,\,U\boxtimes V\mapsto\pi^{*}U\otimes\mathbf{F}(V),$
@@
${\rm Coh}(G/H)_{\bar{g}}\boxtimes{\rm Rep}(H)_{k}\xrightarrow{\simeq}\overline{{\mathscr{C}}_{gH}},\,\,\,U\boxtimes V\mapsto\pi^{*}U\otimes\mathbf{F}(V),$
@@ 7. The center of ${\rm Coh}(G,\omega)$
Throughout this section, we fix $(G,\omega)$ 2.7. Let $\Delta\colon G\to G\times G$ be the diagonal map, $\mathbb{G}:=G\times G$, and $\mathbb{H}:=\Delta(G)$.
Let $\mathscr{Z}(G,\omega):=\mathscr{Z}({\rm Coh}(G,\omega))$ be the center of ${\rm Coh}(G,\omega)$. Recall that objects of $\mathscr{Z}(G,\omega)$ are pairs $(X,{\rm c})$, where $X\in{\rm Coh}(G,\omega)$ and
@@
(7.1) $\mathscr{Z}(G,\omega)\xrightarrow{\simeq}\left({\rm Coh}(G,\omega)\boxtimes{\rm Coh}\left(G,\omega\right)^{{}^{\rm rev}}\right)^{*}_{{\rm Coh}(G,\omega)}$
@@
Let $p_{1},p_{2}:\mathbb{G}\twoheadrightarrow G$ be the obvious projection morphisms, and let
@@
(7.2) ${\rm Coh}(\mathbb{G},\widetilde{\omega})^{*}_{{\mathscr{M}}(\mathbb{H},1)}\xrightarrow{\simeq}\left({\rm Coh}(G,\omega)\boxtimes{\rm Coh}\left(G,\omega\right)^{{}^{\rm rev}}\right)^{*}_{{\rm Coh}(G,\omega)}.$
@@
(7.3) ${\mathscr{C}}\left(\mathbb{G},\widetilde{\omega},\mathbb{H},1\right)\simeq\mathscr{Z}(G,\omega).$
@@
(7.4) ${\rm Coh}^{(G)}(G,\omega)\simeq\mathscr{Z}(G,\omega),$
@@
(7.5) ${\mathscr{C}}\left(\mathbb{G},\widetilde{\omega},\mathbb{H},1\right)\simeq{\rm Coh}^{(G)}(G,\omega).$
@@ 7.1. The structure of ${\mathscr{C}}\left(\mathbb{G},\tilde{\omega},\mathbb{H},1\right)$
    $\mathbf{F}_{{\rm Z}}:{\rm Rep}(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})_{k}\xrightarrow{\simeq}\mathscr{C}_{{\rm Z}},\,\,\,V\mapsto\iota_{{\rm Z}*}\left(\mathscr{O}\left({\rm Z}\right)\otimes_{k}V,\lambda_{V}^{{\rm g}},\rho^{{\rm g}}_{V}\right).$
@@
    $\mathbf{F}_{\mathbb{H}}:{\rm Rep}(\mathbb{H})_{k}\xrightarrow{\simeq}\mathscr{C}_{\mathbb{H}},\,\,\,V\mapsto\iota_{\mathbb{H}*}\left(\mathscr{O}\left(\mathbb{H}\right)\otimes_{k}V,\lambda_{V}^{1},\rho^{1}_{V}\right).$
@@
    There is a bijection between equivalence classes of pairs $({\rm Z},V)$, where ${\rm Z}\in{\rm Y}(k)$ is a closed point with representative ${\rm g}\in{\rm Z}(k)$, and $V\in{\rm Rep}(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})_{k}$ is simple, and simple objects of $\mathscr{C}$}, assigning $({\rm Z},V)$ to $\mathbf{F}_{{\rm Z}}(V)$. Moreover, we have a direct sum decomposition of abelian categories
@@
    For any $V\in{\rm Rep}(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})_{k}$, we have $\mathbf{F}_{{\rm Z}}(V)^{*}\cong\mathbf{F}_{{\rm Z}^{-1}}(V^{*})$.
@@
    For any ${\rm Z}\in{\rm Y}(k)$ with representative ${\rm g}\in{\rm Z}(k)$, and $V$ in ${\rm Rep}(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})_{k}$, we have
@@
    For any ${\rm Z}\in{\rm Y}(k)$ with representative ${\rm g}\in{\rm Z}(k)$, and simple $V\in{\rm Rep}(\mathbb{H}^{{\rm g}},\xi_{{\rm g}}^{-1})_{k}$, we have
@@
[…]
@@ 7.2. The structure of ${\rm Coh}^{(G)}(G,\omega)$
[…]
@@
    $\mathbf{F}_{C}:{\rm Rep}_{k}(G_{C},\omega_{g})\xrightarrow{\simeq}\mathscr{Z}_{C},\,\,\,V\mapsto\iota_{C*}\left(\mathscr{O}(C)\otimes_{k}V,\rho_{V}^{g}\right),$
@@
[…]
    In particular, $\mathbf{F}_{1}:{\rm Rep}_{k}(G)\xrightarrow{\simeq}\mathscr{Z}_{1}\hookrightarrow\mathscr{Z}$ coincides with the canonical embedding of braided tensor categories.
@@ 7.4. $\mathscr{Z}(G,\omega)$ as $G(k)$-equivariantization
Set $\mathscr{D}:={\rm Coh}(G,\omega)$, and $\mathscr{D}^{\circ}:={\rm Coh}(G^{\circ},\omega^{\circ})$. By (2.4), we have equivalences
$\mathscr{D}\simeq\mathscr{D}^{\circ}\rtimes G(k)\simeq\bigoplus_{a\in G(k)}\mathscr{D}^{\circ}\boxtimes a$
of tensor categories, where the associativity constraint on the right hand side category is given by $\omega$ in the obvious way.
@@
Note that $\mathscr{Z}_{\mathscr{D}^{\circ}}(\mathscr{D})\simeq{\rm Coh}^{(G^{\circ})}(G,\omega)$ as tensor categories, and the obvious forgetful tensor functor
@@
(7.11) $F:\mathscr{Z}(G,\omega)\xrightarrow{\simeq}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right)^{G(k)}=\left(\bigoplus_{a\in G(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)\right)^{G(k)}.$
@@
(7.12) $F:\mathscr{Z}(G,\omega)\xrightarrow{\simeq}\bigoplus_{C\in{\rm C}(k)}\mathscr{E}_{C}.$
@@
$F_{C}:\overline{\mathscr{Z}(G,\omega)_{C}}\xrightarrow{\simeq}\mathscr{E}_{C}=\bigoplus_{a\in C(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)^{G(k)}.$
@@
$F_{1}:\overline{{\rm Rep}_{k}(G)}\xrightarrow{\simeq}\mathscr{Z}\left(G^{\circ},\omega^{\circ}\right)^{G(k)}.$
@@
So, let ${\rm V}\in\mathscr{Z}(G,\omega)_{C}$ be simple. By Theorem 7.4, there exists a unique simple $V$ in ${\rm Rep}_{k}(G_{C},\omega_{g})$ such that ${\rm V}={\bf F}_{C}(V)=\mathscr{O}(C)\otimes_{k}V$. It follows that the forgetful image of $F({\rm V})$ in $\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right)$ (7.10) lies in $\bigoplus_{a\in C(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)$, so by the proof of [GNN, Theorem 3.5], $F({\rm V})$ is supported on $C$, as claimed. ∎